help writing the expression in terms of first powers of cosine

Write the expression in terms of first powers of cosine. Do nc 4 sin X= olo cos
Answer #1

sin^4x = (sin^2x)^2

= (1-cos^2x)^2

= 1 + cos^4x - 2cos^2x

= cos^4x - 2cos^2x + 1